A total of 2825 tickets were sold for the concert for a total of $25,898 if advance tickets sold for $8 each and tickets at the door were sold at $10 each, how many tickets of each type were sold?
Found 2 solutions by jorel1380, addingup:
Answer by jorel1380(3719)

Let m be the amount of tickets sold at $10, and n be the amount sold at $8. Then:
m+n=2825
10m+8n=25898
8m+8n=22600
2m=3298
m=1649
n=2825-1649=1176
1649 tickets were sold at the door, and 1176 were sold in advance. ☺☺☺☺

Answer by addingup(3677)

advanced: x
at door: y
---------
x+y = 2,825
thus
y = 2,825-x
:
8x+10y = 25,898 substitute for y
8x+10(2,825-x) = 25,898
8x+28,250-10x = 25,898
-2x = -2352 divide both sides by -2 and remember -/- =
x = 1,176 This is how many advanced tickets were sold. And:
2,825-1,176 = 1,649 tickets were sold at the door